At an output level of 60,000 units, you calculate that the degree of operating leverage is 3.20. Suppose fixed costs are $200,000. What is the operating cash flow at 54,000 units? (Do not round intermediate calculations and round your final answer to the nearest whole number, e.g., 32.)

  

  Operating cash flow $  

  

What is the new degree of operating leverage? (Do not round intermediate calculations and round your final answer to 3 decimal places, e.g., 32.161.)

  

  Degree of operating leverage   
Answer #1

Degree of operating leverage

= Contribution margin / (Contribution margin - Fixed expenses)

Let contribution margin be x

x/(x-200,000) = 3.20

x = 3.20x - 640,000

x = 290,909.09

Contribution margin pu = x/60,000 = 4.84848

Contribution margin for 54000 = 261,818.18

Operating cash flow = (54,000 * 5.84848 - 200,000) = 61,818.18

Degree of operating leverage = 4.24